### v1

# DiLi: A Lock-Free Asynchronously Distributable Linked List

Abstract
Modern databases use dynamic search structures that store a huge amount of data, and often serve them using multi-threaded algorithms to support the ever-increasing throughput needs. When this throughput need exceeds the capacity of the machine hosting the structure, one either needs to replace the underlying hardware (an option that is typically not viable and introduces a long down time) or make the data structure distributed. Static partitioning of the data structure for distribution is not desirable, as it is prone to uneven load distribution over time, and having to change the partitioning scheme later will require downtime.
Since a distributed data structure, inherently, relies on communication support from the network stack and operating systems, we introduce the notion of conditional lock-freedom that extends the notion of lock-free computation with reasonable assumptions about communication between processes. We present DiLi, a conditional lock-free, linearizable, and distributable linked list that can be asynchronously and dynamically (1) partitioned into multiple sublists and (2) load balanced by distributing sublists across multiple machines. DiLi contains primitives for these that also maintain the lock-free property of the underlying search structure that supports find, remove, and insert of a key as the client operations.
Searching for an item in DiLi is by a novel traversal that involves a binary search on the partitioning scheme, and then a linear traversal on a limitable number of linked nodes. As a result, we are able to empirically show that DiLi performs as well as the state-of-the-art lock-free concurrent search structures that are based off of a linked list when executed on a single-machine. We also show that the throughput of DiLi scales linearly with the number of machines that host it.
keywords:
Design and analysis of distributed algorithms, Distributed and concurrent data structures
[…]
Eligible for best student paper award.
1 Introduction
Concurrent dynamic search structures play a crucial role in many modern databases [3, 24, 12]. With the rising number of concurrent operations to a database, memory contention should be managed with care to preserve the atomicity of its operations. For this, locks are often utilized, which can lead to pitfalls such as deadlock, priority inversion, and convoying [18].
Lock-free search structures avoid the drawbacks of lock-based methods but are limited to single-machine shared memory systems. Scaling further will require either a hardware upgrade to the single machine and perform a data migration, or a scheme to distribute load across multiple machines, with each machine handling only a ‘part’ of the data structure. The former can be expensive if the machine is self-hosted (for privacy reasons), or can be limited to the highest available instance if cloud-hosted. Migrations here would still cause downtime to the database. Further, the need to distribute also arises from the observation that Moore’s law now has diminishing returns [22].
Partitioning for load balancing can be static (fixed key distribution per machine) or dynamic (key distribution evolves over time). Static schemes cannot adapt to workload shifts or scale without downtime. Dynamic schemes offer flexibility, but require shared memory structures to be temporarily unavailable during redistribution. It is, therefore, necessary to develop dynamic partitioning that supports redistribution asynchronously, i.e., while preserving the properties of the client operations.
In this paper, we focus on conditional lock-free linked lists. Linked lists are fundamental data structures that can in turn construct other data structures [30]. Specifically, we distribute the well-known Harris Linked list [14], and improve its performance even on a single-machine setup. Due to the distributed scalability and the general performance improvement, any multi-threaded lock-free data structure that is implemented with Harris List as its foundation, can automatically become a faster and distributable data structure by replacing Harris List with DiLi for its foundation.
Contributions: We introduce a new distributed linked list DiLi – that performs find, insert and remove of a key both in a multi-threaded and a multi-machine setup while supporting asynchronous dynamic partitioning. Any part of this data structure (henceforth called a sublist) can be split into two smaller parts. The resulting partitioning scheme additionally serves as an index for quick search traversals, resulting in a significant performance boost that is sufficient to become comparable to lock-free skip lists [11] in workloads of varying read proportions (as we show empirically).
By splitting the data structure into sublists, each sublist can be individually moved to another machine, thus switching the load of this sublist to the other machine. We refer to this as a dynamic load distribution. The dynamic load distribution is challenging to implement in an asynchronous manner. We provide an interface to implement any dynamic load balancing scheme using our “Split” and “Move” operations.
To achieve the above, we only additionally require a single-word compare-and-swap (CAS) and fetch-and-add atomic instructions, which are now available in most commodity hardware.
Empirically, we show that 1. The performance of DiLi is comparable to that of a lock-free skip list in multiple workloads, 2. DiLi is capable of automatic scaling upto several machines with a linear growth in throughput, and 3. The performance of Split and Move operations are practical even under write-intensive workloads.
Outline: We introduce the notion of conditional lock-freedom in Section˜2, introduce our system model in Section˜3, and explain our design principles in Section˜4. DiLi is then described through the following three sections: Algorithm (Section˜5), Correctness (Section˜6) and Empirical Evaluation (Section˜7). Then we compare DiLi with related work in Section˜8. Finally, the concluding remarks are in Section˜9.
2 Extending Lock-Freedom for Distributed Systems
Even in simple client-server models such as a single server-client model, the traditional lock-freedom definition: “A method is lock-free if it guarantees that infinitely often some method call finishes in a finite number of steps” [18] is difficult to achieve, as any client request has to get past unbounded network delays to get the operation request on a data structure served. Also, the client thread that initiates this request typically goes to sleep (hence having a dependency on another thread) until it receives a response from an associated server machine thread. In other words, the definition of traditional lock freedom cannot be satisfied in any client-server machine (or more generally in distributed systems). Due to the above constraints, we condition any communication between threads (such as messages) in the distributed system to only take a finite number of steps (e.g., with reliable channels).
Definition 2.1 (Conditional Lock-Freedom).
We say that a method (or operation) implementation is conditional lock-free if it is traditional lock-free when the execution satisfies the following condition: any communication between threads (such as messages) in the distributed system only takes a finite number of steps (e.g., with reliable channels).
We have chosen to define conditional lock-freedom in this fashion to distinguish between a simple scenario where a server in a client-server model (1) implements the Harris list, and (2) implements a traditional lock based list. The former will be conditional lock-free while the latter would not.
3 System Model
We study a list sorted by key and refer to each of its elements as elements. To construct our partitioned version of this list, we introduce the notion of a sublist that comprises a set of items contained by a subhead and a subtail. For example, in Figure˜1, we show a list consisting of 3 sublists (SL1 to SL3), each with a subhead (SH) and a subtail (ST) that hold a range of keys. The list supports the following three operations: 1. Find(key): Returns false if the key is not found in the list, and otherwise returns true. 2. Insert(key): Returns false if the key is already in the list and otherwise inserts the key in its sorted position, returning true. 3. Remove(key): Returns false if the key is not found in the list, and otherwise deletes the key, de-links the item off the list, and returns true. Additionally, a server may invoke operations Split (to partition the submit into two sublists), Move (to create a live clone of the sublist to another server), and Switch (to transfer ownership of sublist to new server). [^1] We also support a Merge operation for two adjacent sublists. Due to the similarity of its logic to the Split algorithm, we have pushed it to Appendix B.
Figure 1: A sample list for keys 1–100 being composed of a set of 3 sublists.
We consider a client-server model of a distributed system with any number of multi-core machines acting as servers. The server and client machines are fully connected by reliable channels (finite time to deliver a message) that guarantee eventual delivery. All messages exchanged in the channels are via remote procedure calls (RPCs). A server may not invoke any RPC on the clients. However, servers may invoke RPCs on each other and also send messages to each other asynchronously.
As shown in Figure˜2, we follow a decentralized architecture. Each client will be assigned a specific server X (X is any server in the system) to invoke its list operations for any key $k$ through RPCs. The server X responds to the client, by invoking RPC to server Y that is capable of performing the list operation for $k$. In this case, X acted as a proxy and is said to have delegateddelegated the request to Y. Note that X could also receive delegated requests from other servers as it also owns some sublists. Naturally, in a single machine setup, this overhead is avoided, as X will always be the same as Y. Further, we assume that a server machine does not crash entirely, and has atleast one usable core. This is essential to serve the keyset that the machine is responsible for, as we do not create replicas of the list.
An important concern when aiming for linearizability in a distributed system is to operate under the constraints of the CAP theorem [7]. By ensuring that any operation takes place on only one sublist, and that a sublist entirely resides in one single machine at any given time, we are able to overcome the effects of network partitioning that impedes linearizability.
We allow a sublist to undergo a Split or Move or Switch by only one thread at a time. This can be easily achieved by assigning only one background thread per machine to perform such operations. Note that even if we have atmost one background thread operating on a sublist at any time, we are required to ensure the lock-free nature of the client operations on the list. We refer to this guarantee as the asynchronous nature of the background operations.
Figure 2: Client invoking an RPC request on a key to its assigned Server X and server X delegating the request to server Y that owns the sublist for that key.
4 Design Principles
The design of the distributed list involves several techniques in order to provide a linearizable lock-free implementation. The first is a smart pointer, which is inspired by the Harris Linked List [14]. Harris used the least significant bit (LSB) of a list item’s next pointer to denote whether an item has been soft deleted, as it is otherwise always set to 0 due to address alignment. In doing so, the time point of checking the value of the next pointer became crucial for linearizability arguments. In addition, there is common limitation on the x86-64 MMUs of commodity hardware (such as without 5-level paging extensions [1]) only to have a 48-bit virtual address space. Due to this, the 16 most significant bits of the 64-bit pointer remain unused during memory allocations. DiLi uses these 16 bits to globally indicate the server ID that the distributed list item belongs to.
Secondly, we maintain all sublists as entries in a lazily replicated registry—an index where each entry has a smart subhead pointer that encodes the server ID. This ID identifies the server that hosts the sublist and guides request routing. Although lazy replication may cause temporary staleness and latency, safeguards (see Section˜5) ensure that lock-freedom and linearizability remain intact. Lazy replication allows the registry to remain sorted via copy-on-write inserts. This enables logarithmic-time key lookup. This leads to our next design principle – the hybrid search where we use a logarithmic search in the registry followed by a linear search in the sublist. By keeping a check on the number of items present in a sublist, we achieve very quick traversals in the distributed list.
The Move operation handles the load balancing of the distributed list. The Move creates a momentary live clone of the sublist in the new machine. To achieve this in a non-blocking manner [^2] Note that it is insufficient to use snapshotting algorithms for this purpose, because we need a snapshot that always matches the live state of the sublist, instead of a state at some point of time in the past., we perform temporary replication of updates that happen during the move. For instance, if insert($k$), remove($k$), and insert($k$) occur in order, both replicas must retain $k$ - avoiding out-of-order replay like insert($k$), insert($k$), remove($k$). This challenge stems from mismatches between the linearized update order and the delivery order. Our novel replay algorithm achieves this by tracking a logical timestamp on the final location of the item.
The move terminates, activates the cloned sublist, and transfers ownership to the new server when both machines have identical copies of the sublist. Although blocking client operations simplifies this, our approach achieves it asynchronously using integer counters around sublist updates. The Switch operation then needs to update the registry in all servers before another Move can happen on this sublist. In the interim, careful rerouting (called delegations) of requests should take place as the registry can be temporarily out-dated.
5 DiLi: A Distributable List
DiLi is a linearizable sorted linked list with a two-layered data structure (described in Section˜5.1) – (1) an index called the Registry, that book keeps the subhead and subtail pointers (to shortcut into a sublist), and (2) sublists that consist of the underlying list items. Upon reaching a subhead, a linear traversal is employed to find the search node. The find, insert, and remove operations (as described in Section˜5.2) that utilize this traversal are lock-free, but only operate on the sublist space. For ease of explanation, we use colors in the algorithms. Black lines are executed the same way, regardless of whether there is an ongoing background operation, the red lines adapt to a concurrent Split and hence will be explained as part of Section˜5.3, and the blue lines adapt to a concurrent Move/Switch and are explained in Section˜5.4. The novel contribution in our approach is an asynchronous yet versatile Split, Move, and Switch algorithms that can be used to build any partitioning and load balancing scheme of choice to suit the workload of interest.
5.1 Data Structures: List Item and Registry
Algorithm˜1 depict the two building blocks of DiLi. Like any list item, DiLi items contain a key to search for the item and a next pointer to lead to the next item. Note that this is a distributed list, with a traversal guided by the next pointer. We denote the type of this pointer as Ref. Any Ref pointer, say X, needs to be operated on by a bit mask to force the first 16 bits and the last bit to 0 before accessing the address that X points to. We overload this functionality into the C++ pointer access notation ($\rightarrow$). Fetching serverID bits, and if the item has been marked, are also implemented through bitmasks. For convenience, we denote them as X.id and X.mark in the algorithms.
The registry consists of sublist entries and supports getByKey(key) operation to find the sublist entry to which a searched key belongs. Operation addEntry(entry) is used to insert an entry into an already sorted registry. We utilize a copy-on-write (COW) technique to update the registry pointer, to atomically update the registry with a new sorted registry with an added entry. Because registry replication is done lazily, it is sufficient to have only one update happening at a time on the registry. Hence any balanced single-writer tree data structure (binary tree, B-tree, or even a skip list) could also be used to build the registry, thereby offering a logarithmic search for getByKey(key) operation. For the sake of simplicity, we maintain the registry as a binary search tree in a large array and keep track of its size. We also utilize a variant of hazard pointers [25] to reclaim memory when an old pointer of the registry (after overwritten by COW) is no longer referenced by any thread. To save space, the functioning of the registry is explained in Appendix˜A.
1
2struct Item:
3  Key key;
4  Key keyMax;
5  int ts;
6  int sId;
7  Ref next;
8  int stCt ;
9  int endCt ;
10  Ref newLoc ;
11
12struct Entry:
13  Ref subhead;
14  Ref subtail;
15  Key keyMin;
16  Key keyMax;
17  int stCt ;
18  int endCt ;
19  int offset;
20
21struct Registry:
22  Entry * entries[MAX_SUBLISTS] = $\{null\cdots null\}$;
23  int size = 0;
24
25
Algorithm 1 Data Structures: Item and Registry
5.2 Client Operations
Here, we present the functioning of Find, Remove, and Insert operations assuming that the partitioning index (registry) and the ownership of various sublists by different machines stay static. When that assumption breaks, the corresponding changes to their execution are explained in Sections 5.3 (red lines) and 5.4 (blue lines), respectively.
All client operations perform the following in order: 1. A binary search on the replicated registry to retrieve the sublist entry for a key. 2. Routing the request to the appropriate server based on the server ID of the subhead present in the sublist entry. 3. A linear traversal on the sublist to perform operations in a Harris-List fashion with retries beginning from the subhead of the sublist.
A registry look-up is first performed for any operation requested by the client. This returns the subhead entry for the key range, that in-turn contains the $\tt Ref$} for its subhead. By checking the id bits of Ref, the machine serving the sublist is determined to reroute as needed. The operations take only the key as a parameter if the request comes directly from a client, and if it is instead a delegation, the subhead is also a parameter. A re-routed request already has the subhead required, and hence skips the registry lookup step on its machine.
The end machine that serves the sublist traverses using the Search algorithm in Algorithm˜2. It resembles Harris List search, where the left and right nodes to the searched key are returned, while delinking any marked items encountered along the traversal. However, we optimize this search by limiting the search to start from the pre-determined subhead, instead of the list head. In our case, the rightNode is the item of the searched key or the one immediately greater present in the list. leftNode is the corresponding previous item.
Find and Remove utilize the returned right node of a search to determine if a key is present. Find returns its response based on this check (after Line 2), while Remove continues with a Delete on the found item using the Delete function. The Delete attempts to soft delete the node by marking its next pointer’s mark bit using CAS. If the CAS succeeds (Line 2), then the Remove returns true, and if it is already marked, then it returns false.
The Insert operation as described in Algorithm˜3 utilizes the InsertInSublist routine after determining the sublist to insert the key. This routine begins with a search operation to determine the left and right node window. If the right node matches the key to be inserted, then the Insert returns false. If the next pointer of left node is altered by getting marked or through a competing insert, then the CAS fails and the search is restarted on the same sublist to retry the insert. The logic behind the stCt, endCt, newLoc and ts assignments become apparent only during the background operations, as covered in the next sections.
26
27Ref Search(Key key, Ref head, Ref &leftNode):
28  curr = head;
29
[…]
32
33  while true do
34     leftNode = prev = curr;
35     $curr\rightarrownext$;
[…]
38        if head.id == me
39           leftNode = prev = head;
[…]
43
44          continue;
45
46       return head;
47
[…]
50           return null;
51
52        $curr\rightarrownext$;
53        if curr.id $\neq$ me
54           return curr;
55
56       leftNode = prev = curr;
57        head = curr;
58        $curr\rightarrownext$;
59
[…]
61        if delinkNode(prev, curr)
62           Lines 2 to 2;
63           continue;
64
65       else
66           leftNode == prev == head;
[…]
70
71           continue;
72
[…]
74        return curr;
75
[…]
77        return null;
78
[…]
81
82  return null;
83
84bool delinkNode(Ref prev, Ref &curr):
[…]
86  while currNext.mark = 1 do
[…]
88
89 temp = curr;
90  if CAS(prev,temp,currNext)
91     curr = currNext;
92     return true;
93
94 return false;
95
96bool Find(Key key, Ref SH = null):
[…]
98     entry = registry.getByKey(key);
[…]
100
101 sId = SH.id;
102  if sId == me
103     leftNode = null;
104     node = Search(key, SH, leftNode);
105     if node == null
106        return false;
107
108    sId = node.id;
109     if sId == me
110        return true;
111
112     response = Send Find(key,node) to sId;
113     return response;
114
115 response = Send Find(key,SH) to sId;
116  return response;
117
118bool Remove(Key key, Ref SH = null):
119  Same as Find but line 2 is changed to
120 return Delete(node, key);
121  and the corresponding delegates are that of Remove instead of Find.
122bool Delete(Ref node, Key key):
123  result = false;
[…]
125     return false;
126
[…]
128
[…]
131     return response;
132
133  while true do
[…]
136        break;
137
[…]
139     newVal = temp;
140     newVal.mark = true;
[…]
142        result = true;
[…]
145
146       else
[…]
148
149       break;
150
151 leftNode = null;
[…]
153  Search(key, SH, leftNode);
154  return result;
155
Algorithm 2 Distributed List Traversal, Find and Remove
156
bool Insert(Key key, Ref SH = null):
[…]
158     entry = registry.getByKey(key);
[…]
160
161 sId = SH.id;
162  if sId == me
163     return InsertInSublist(key, SH);
164
165 response = Send Insert(key,SH) to sId;
166  return response;
167
168Entry *Split(Entry * entry, Ref sItem):
169  *newstCt = new int[1];
170  *newendCt = new int[1];
171  SH = new Item(SH_KEY, $\phi$, ts.fetch_add(1), me, null, newstCt, newendCt, null);
[…]
173
174 do
[…]
176     if temp.mark
177        return null;
178
[…]
181
[…]
184  do
185     prev = curr;
[…]
188     $curr\rightarrownext$;
189
[…]
191 do
192
193    $a_{1}$ = newstCt- newendCt ;
194     $a_{2}$ = $\rightarrowstCt- sItem\rightarrowendCt$ ;
195
196 while $a_{1}+a_{2}\neq entry\rightarrow offset$ ;
[…]
198
199 registry.addEntry(newEntry);
[…]
203
204 for $i\in(serverList-\{me\})$ do
[…]
206
207 return newEntry;
208
209bool RegisterSublistRecv(Key keyMin, Ref SH):
210  Entry leftEntry = registry.getByKey(keyMin);
[…]
213  return true;
214
215bool InsertInSublist(Key key, Ref SH):
216  while true do
217     leftNode = null;
218     rightNode = Search(key, SH, leftNode);
219     if rightNode $\neq$ null
220        if rightNode.id == me
221           return false;
222
223       response = Send Insert(key,rightNode) to rightNode.id;
224        return response;
225
[…]
229           return false;
230
[…]
232        if SH.id == me
233           continue;
234
[…]
236        return response;
237
[…]
241           continue;
242
[…]
245        Lines 3 to 3.
[…]
250
251       else
[…]
253
254       return true;
255
[…]
257
Algorithm 3 Operations: Insert and Split
5.3 The Split Algorithm
The Split operation (Algorithm˜3) takes two inputs – a sublist entry and an item (referred to as sItem) at which the sublist is to be split. This truncates the original sublist to a smaller key range and creates a new sublist for the remaining key range. The Split is performed in through 4 major steps, while always guaranteeing that the client operations stay lock-free:
  1. (1)
    Allocate new counters stCt and endCt (lines 3 and 3) for the new sublist. The purpose of these counts is to ensure that if a client operation overlaps with Split/Move, it is detected so that Split/Move can take corrective action.
  2. (2)
    Insert a connected block of a subtail followed by a subhead (as in Lines 3 and 3) at the next pointer of sItem (insertion is in Line 3). If sItem gets deleted before this, split operation fails (Line 3) and can be retried with another call at a different sItem.
  3. (3)
    Update the stCt and endCt pointers of the second half of the sublist (Lines 3 to 3). Note that if an item gets inserted and gets missed in the traversal, it must have copied the counters from the item it was previous to, whose counters (or their ancestors’) were updated during this traversal. Thus, they must also contain the new counter pointers.
  4. (4)
    Update the registry by adding an entry for the second half of the sublist, and then truncating the first half to its reduced keyMax (Lines 3 to 3).
The above approach has two corner cases to handle due to the asynchronous insertion. First, for client operations that concurrently search for a key higher than sItem, the search can get truncated at the inserted subtail in case the client operation traversal encounters it. For this purpose, the subtails contain a keyMax element stored. Specifically, the red lines from 2 to 2 check whether the key belongs to the first half of the sublist that is undergoing a concurrent split or to the second half. If it is in the second half, then its subhead pointer is updated, speeding up potential retries of the lock-free Harris List traversal.
Secondly, we can no longer expect the difference between stCt and endCt of the individual sublists to be 0 when there are no ongoing updates (insert/remove). This is because, during the counter updates to the second half of the sublist, some updates could have incremented the stCt of the left half and updated the endCt of the right half. This accumulated change to the counter difference when the sublist is static is denoted as the offset of the sublist. As would be explained later, the new offset is important for Move and Switch operations. To compute this new offset, we employ a spin (Lines 3 to 3) that breaks only when the computed values $a_{1}$ and $a_{2}$ reflect the new right half offset and left half offsets respectively. The condition works because the difference between the stCt and endCt of any sublist stabilizes to a minimum only when there are no ongoing updates in the sublist. This is true for both the left and right halves of the sublist, and since the increments went to exactly one of the two stCt counters and one of the two endCt counters for each update, the sum of the new offsets must be equal to the old offset when the two new sublists together have no ongoing updates. We note that this spin does not delay the client operations, and only delays background operations on the split sublists.
5.4 Move, Replay and Switch Algorithms
A Move routine invoked on a server machine (henceforth $S_{1}$) takes the registry entry of a sublist and the machine to move it to (henceforth $S_{2}$), as parameters. The goal of Move is to create a clone of a sublist and then make it the live copy. To achieve this, the Move routine as described in Algorithm˜4 traverses each item (by recursively calling the MoveItem routine) in the sublist in $S_{1}$ and creates a copy of them in $S_{2}$. The reference in $S_{2}$ of last moved item is sent back each time so that subsequent items can be linked to the previously moved item.
Move compensates for concurrent updates that the Move routine did not witness during its traversal, by 1. detecting that such concurrent updates occurred, and 2. replicatingthose updates correctly on $S_{2}$. The former is achieved by using stCt (incremented before the start of update operation) and endCt (incremented after the update operation); and the latter is achieved by sending replicates (Line 4 in Algorithm˜2 and Line 3 in Algorithm˜3) of these operations to $S_{2}$. As discussed in Section˜4, the inserts and removes will need further specification for the correct replay of the updates in $S_{2}$, as they may appear out of order when received. To account for this, the replay models the changes to the sublist as a result of the following update operations: InsertAfter(Ref, Item) and Delete(Ref), that distinctly convey where the insert CAS (Line 3) and remove CAS (Line 2) that had succeeded. If multiple items got inserted at a particular Ref $A$, that had been moved by the Move routine, then the item with the lower timestamp got inserted first. As for the Delete, the same item that was moved to $S_{2}$ is the one that gets deleted, and a subsequent insertion for the same key is treated as a different item. Since Ref present in $S_{1}$ for an item is not known by Ref copy of it in $S_{2}$, we denote the uniqueness of an item by the <sId, timestamp> tuple.
Using the above observations, we developed our Replay algorithm presented in Lines 4 to 4. The received move of an item and a received insert replicate utilize this routine. Once the prev item is identified using their <<sId, timestamp>> tuple, the Replay algorithm takes this prev item and traverses from prev to find the right position to insert the item. To not miss any item, the Move routine even moves marked items as part of the algorithm, which are then delinked once the cloned sublist becomes active. Because replay is concurrently invoked by Move and multiple Replicate threads, insertions at replay are retried. An extensive proof of correctness for the replay is provided in Section˜D.3.
Finally, to detect that a sublist is live in $S_{1}$, we utilize the stCt and endCt values of the sublist. All updates increment the stCt before performing the update, and increment the endCt after performing the update. However, if a replicate message is sent, the endCt is incremented only after the completion of the replay (Line 4 and 4). This way, when the endCt measured matches the stCt measured immediately after, we can be sure that both copies of the sublist are at the same state. At this point we use a CAS to set the stCt to -$\infty$ (Line 4). All updates check if the stCt is negative after they increment it. If it is negative, they delegate the request to $S_{2}$. Otherwise, they continue with the operation, delaying the Move until the update increments endCt and so never block client operations.
After stCt is permanently negative (the blue lines other than send replicates reflect this scenario), all updates are redirected towards $S_{2}$, thereby making its sublist version live and leaving the sublist at $S_{1}$ stale. This point begins the Switch operation (or phase). The Switch routine in Algorithm˜5 is invoked to first complete the dynamic re-distribution by updating the next pointer of the previous subtail. Since the current Switch has no control over the background operations on the sublist previous to it, this attempt is done in a loop to account for redirections (expected to not happen more than once, as it is highly undesirable to move the same sublist way too frequently). After this step, the Switch, and hence the re-distribution is completed by updating the registry entry for the sublist in all servers.
258
Move(Entry entry, int newSId):
[…]
260  remoteSH = Send MoveSH($\rightarrowsId, head\rightarrowts$, $\rightarrowkeyMax, entry\rightarrow$);
[…]
263  do
264     temp = $\rightarrowendCt + entry\rightarrowoffset$;
265
[…]
267 Switch(entry, newSId);
268
269MoveNext(Ref prev, Ref curr, int newSiD):
[…]
273
[…]
275     Send RepDelete($\rightarrownewLoc, curr$, $\rightarrowts, curr\rightarrowsId$) to newSId;
276
[…]
278     return;
279
280 MoveItem($\rightarrownewLoc, curr\rightarrownext$, newSId);
281
282Ref MoveSHRecv(int itemSId, int itemTs, Key keyMax):
283  *newstCt = new int[1];
284  *newendCt = new int[1];
285  ST = new Item(ST_KEY, keyMax, , , null, newstCt, newendCt, null);
286  SH = new Item(SH_KEY, $\phi$, itemTs, itemSId, ST, newstCt, newendCt, null);
287  entry = registry.getByKey(keyMax);
[…]
292  return SH;
293
[…]
295  while $\rightarrowts \neq prev$Ts $\lor$ $\rightarrowsId \neq prev$SId do
296     $prev\rightarrownext$;
[…]
298        prev = prevNewLoc;
299
300
301 return <<prevNewLoc, Replay(prev, itemTs, key, itemSId, itemTs, false)>>;
302
303Ref RepDeleteRecv(Ref prevNewLoc, int prevSId, int prevTs):
304  Lines 4 to 4;
305  do
[…]
307     ref = temp;
308     ref.mark = true;
309
[…]
311 return prevNewLoc;
312
313Ref MoveItemRecv(Ref prev, Key key, bool isMarked, Ref STNext, int itemSid, int itemTs):
314  if key == ST_KEY
315     curr = prev;
316     do
317        $curr\rightarrownext$;
318
[…]
321     return curr;
322
323 return Replay($prev\rightarrowts,$ key, itemSId, itemTs, isMarked);
324
325Ref Replay(Ref prev, Ref compTs, Key key, int itemSId, int itemTs, bool isMarked):
326  do
327     curr = prev;
328     do
329        currPrev = curr;
330        $curr\rightarrownext$;
331
[…]
336     ref = newItem;
337     ref.mark = temp.mark;
338
[…]
340 return newItem;
341
342InsertReplayResponseRecv(Ref oldLoc, Ref newLoc):
[…]
345
346RemoveReplayResponseRecv(Ref oldLoc):
[…]
348
Algorithm 4 The Move and Replay Algorithm
349
Switch(Entry entry, int newSId):
[…]
352     leftEntry = registry.getByKey(keyMin);
[…]
354     do
355        if leftSH.id == me
[…]
357
358       else
[…]
360
361       if leftSH == null
362           response = true;
363
364    while $\neg$ response;
[…]
366
367 for $i\in serverList-\{me,newSId\}$ do
368     response = Send SwitchServer( keyMax, newSH) to i;
369
370 response = Send SwitchServer(keyMax, newSH) to newSId;
371
372SwitchServerRecv(Key keyMax, Ref newSH:
373  entry = registry.getByKey(keyMax);
[…]
375  return true;
376
377Ref SwitchSTRecv(Key keyMin, Ref newSH):
378  leftEntry = registry.getByKey(keyMin);
[…]
380  if leftSH.id == me
[…]
382        return null;
383
[…]
385
386 return leftSH;
387
388bool switchNextST(Ref leftST, Ref newSH):
[…]
391     return false;
392
[…]
395  return true;
396
397
Algorithm 5 Switch Algorithm
6 Correctness
We show the correctness of our algorithms by showing that the client operations are linearizable, that they never get blocked by the background operations. We push the correctness of the background operations to Appendix˜D to save space. However, for the rest of this section, we borrow the following observations proved in the appendix: 1. There is only one active subhead to a sublist at any given time. 2. If an update on a sublist has checked that the sign of stCt is positive after incrementing it, then the stCt of the sublist cannot become negative until the update also increments its endCt.
6.1 Linearizability
We show that the client operations are linearizable by identifying linearization points [19] (LP) for every possible output. DiLi is designed in a way that every operation can be said to take place in some specific machine entirely, by identifying the sublist that the operation takes place in. This is done by noting the currently active subhead for the sublist at the time when the operation is said to be linearized. Thus, our proof sketch for linearizability is by identifying the same linearization points as in the Harris List, while also showing that the same point in time was on the machine with the active sublist.
Find: For a successful Find, the LP is immediately after the execution of the mark check returned false at Line 2 at the machine where Line 2 also had its start count to be non-negative. For an unsuccessful Find, we have two cases – 1. If Line 2 ever was evaluated as true, the LP is after the execution of the key comparison in that line and it took place in the machine where the start count check on the same line returned false. 2. If the above was not satisfied, then a subtail that upper bounds the key must have been reached by the check on Line 2. Since a subtail does not get updated, instead of performing a subsequent start count check here, we instead linearize the find after the execution of Line 2 to be false and the operation took place in the machine where the latter was evaluated to be false.
Remove: For a successful Remove, the LP is immediately after the successful execution of CAS in Line 2, where the next pointer of the item of the matching key gets marked. For an unsuccessful Remove, we have two cases – 1. If the search returned null, The LP is the same as an unsuccessful Find. 2. If the search returned an item of a matching key, the LP is immediately after Line 2 returned true, i.e., when the node is already seen as marked/removed by another concurrent remove of the same key.
Insert: For a successful Insert, the LP is immediately after the successful execution of CAS in Line 3. For an unsuccessful Insert, we similarly have two cases – 1. If the search returned null, the LP is the same as an unsuccessful Find. 2. If the search returned a rightNode of the matching key, the LP is the same as a successful Find.
6.2 Conditional lock-freedom of Client Operations
Conditional lock-freedom depends on the fact that any client operation utilizes atmost four threads (three when there is no maintenance) and that the number of instructions executed by a delegating thread is bounded (proved in Section˜D.2). When there is no background operation, being similar to a Harris List, a successful Remove causes at most two updates to a sublist – one for marking (by CAS in Line 2) and one for delinking (by CAS in Line 2). A successful insert causes at most one update (by CAS in Line 3). The number of retries in Search is limited by the number of marked nodes to be delinked from concurrent Removes. Hence, no operation requires more retries than the number of concurrent updates [14].
During a Split operation, the search traversal (as denoted by the red lines in Algorithm˜2) is only enhanced by an earlier truncation from the introduced subtail. An earlier truncation that occurs exactly at the item where a concurrent insert is taking place shortcuts to the next subhead (red lines of Algorithm˜3, as the insertion is now expected to happen on the subhead instead of the item where the split is taking place, in order to preserve the key range property of the split sublists. By having zero additional search traversals for find and remove, and exactly one additional search traversal, the Harris List based operations that were lock-free when there is no split operation, are also lock-free during a Split operation.
During a Move operation (before Switch), the Find operation functions in the usual Harris List manner. The Remove and Insert operations additionally send Replication messages (replicates are asynchronous messages for which we do not wait for a response). Thus, Remove and Insert on a sublist only perform a constant amount of extra steps due to a concurrent Move on the same sublist.
During a Switch of a sublist, until the sublist entries of all servers are updated with the latest subhead, there will be an additional delegation to the right machine using the blue lines of Algorithm˜2 and Algorithm˜3 where the stCt values are checked to be negative. Thus, any concurrent client operation on the sublist during this time suffers from exactly one additional network hop and search traversal on the sublist. Since the search traversal and client operations were lock-free in the absence of a Switch, the client operations are lock-free even during the Switch.
7 Empirical Evaluation
7.1 Implementation
We implemented DiLi in C++11 in the Ubuntu 24.04.1 LTS operating system and utilized Google’s RPC framework (gRPC) v1.66 [13]. The RPC requests are processed asynchronously based on thread availability, while the requester of any RPC synchronously waits for a response. The only situation a requester instead requests asynchronously is to send replication of updates whose responses are processed as asynchronous callbacks by a separate group of threads.
Each machine that serves DiLi is assigned an initial key range to serve the list, chosen naively by a range partitioning on the key range of the list. We also implement a naive load balancer that utilizes the Split, Move, and Switch algorithms. The load balancer is a separate thread spawned in each machine to repeatedly traverse through all sublists held by the machine and to find sublists that are bigger than a threshold of 125 in size, and to use Split roughly in the middle to create smaller sublists for quicker traversal. Additionally, a rough size of the list is measured by taking the individual sum of the sublist sizes and exchanged between machines to make a load balancing decision. A decision to move is made when a machine holds more than 110% of its assigned load, and invokes Move on one of its sublists to a machine with the least load. In this way, the Split invocations control the amount of linear traversal on the sublist for client operations, and the Move invocations roughly balance the number of items present in each machine. Using the same gRPC framework, we also implemented the standard lock-free skip list [11] in C++, and the Harris Linked list for comparison. All implementations can be found in this anonymized repository [23].
7.2 Setup and Methodology
We perform two main experiments to evaluate the performance of DiLi: 1. A single-machine comparison of DiLi with a lock-free skip list and the Harris Linked list. 2. A Distributed Scalability test to show that DiLi is scalable as we add more number of machines.
Both experiments utilize Yahoo! Cloud Serving Benchmark (YCSB) workloads [8] to simulate workloads of a Zipfian distribution. We do a load of 1M keys into the list and then perform the experiment for a workload of 2M operations, for three different workload read-write proportions : 10% read, 50% read and 90% read. The write proportion is equally split between inserts and removes, to keep the size of the list roughly the same.
The experiments were performed using C7i instances from Amazon Web Services (AWS) as the server machines for the list. The instance features an Intel(R) Xeon(R) Platinum 8488C (2.4GHz, 105MB L3 cache), with each core supporting 2 way hyperthreading. The single-machine comparison utilized an 8 core instance of it with 32 GB of RAM, and the distributed list comparison had multiple 2 core instances, each with 8 GB of RAM.
The results include the throughput measured as the number of operations completed per second. To obtain the maximum throughput achieved by these algorithms, we perform several executions of the same experiment, varying the number of clients through a separate large compute instance, until the server machines reach close to the maximum CPU utilization ($\sim 100\%$ as monitored on AWS CloudWatch). A dry run of 2 iterations is first performed before running either experiment to warm up the servers. Then each experiment is repeated 3 times and the average of those executions is used to compute the reported results.
7.3 Results and Discussion
(a) Single Machine Performance of DiLi as compared to skip lists and Harris List.
(b) Distributed Scalability of DiLi. Here DiLi_x stands for DiLi with x number of similar servers.
Single-Machine Comparisons: The results are in Figure˜3(a). As expected, the use of Split in the background to generate smaller sublists permits significantly faster search compared to Harris List. In fact, DiLi provides a comparable performance to skip lists, that are designed to permit faster traversals on a list. A key parameter for skip lists is the maximum number of levels that it permits. For 1M items, 25 levels are sufficient to have a $\log(n)$ expected traversal time. Hence we test it for multiple levels (10, 25 and 100) around it. From this result, we observe that the ability to horizontally scale across multiple machines has not created a loss in throughput if DiLi is run on a single machine. DiLi provides additional benefits over a skip list. Specifically, the memory usage after the 1M item load was around 370 MB for the 25-level skip list, whereas in DiLi it was only 170 MB. This is expected because the memory usage of a skip list grows by an additional factor of the number of levels in the skip list. We do not provide client operation latency results here, as we observed that the majority of the latency for skip lists and DiLi were caused only by the round-trip delay between the client and the server.
Distributed Scalability: We evaluate the scalability of DiLi by comparing the maximum throughput with 2, 4, 6 and 8 servers (cf. Figure˜3(b)) for each of the aforementioned workloads. Interestingly, we not only observe our claim that DiLi scales linearly for every workload, but we also observe that the scaling differs based on the type of workload. This illustrates how our naive implementation of the load balancer affects the rate of growth in performance. This calls for future work on how the Split and Move interface could be properly utilized to optimize performance for a chosen workload.
Performance of Background Operations: We additionally performed a minor experiment in a 2 machine setup, where the first machine serves the entire key range for the list, and a second machine starts without owning any part of the key range. We performed the 1M key load onto this setup, and set the load balancer to transfer sublists to the idle server dynamically during an insert load, until both machines have comparable number of items. During this test, we measured the latency of the Split and Move operations (as depicted in Appendix˜C). The average latency to Split was 2.65 ms and the average latency to Move a sublist was 51.53 ms. Thus, we empirically verify that the asynchronous operations can terminate quickly even during an insert dominated workload.
8 Related Work
Linked lists are fundamental data structures. A sequential linked list is easy to implement. However, a concurrent link list is challenging to design, and hence has been studied extensively for decades. Consequently, there exist many lock-free implementations – [30, 14, 9]; and a few wait-free implementations – [31, 29]. The first lock-free implementation using the atomic CAS primitive (for a non-distributed list) is due to Valois [30]. Harris [14] gives a lock-free ordered list implementation. Among lock-free lists with no indexing and versioning, Harris List has been the state-of-the-art. Heller et al. [16] give a lock-based linked list design with a wait-free lookup operation. Our linear traversal of client operations resembles that of a Harris list. A wait-free ordered linked list implementation is given by Timnat et al. [29].
Some papers focused on optimizing the cost of operations in addition to correctness and progress guarantees established in the aforementioned papers. An example is a lock-free list implementation due to Fomitchev and Ruppert [9], with worst-case linear amortized cost. [14] and [9] both utilize a $\tt{HelpMarked}$ in their read operations. They perform de-linking of marked nodes during any traversal. Zhang et al.[31] provided an unordered linked list where insertions always occur in the head position of the linked list.
A concurrent lock-free doubly-linked list implementation is given by Attiya and Hillel [4], using a double CAS operation. This doubly-linked list does not have a read operation. The same is implemented using only CAS by Sundell and Tsigas [27].
The aforementioned implementations typically consider the sequences of entries in the linked list to reside in different chunks of memory. To take advantage of entries that are instead on the same chunk, Braginsky and Petrank [6] developed a concurrent linked list that maintains in each memory chunk, a certain number of entries within a defined minimum and maximum. Their approach shares some similarities with our key range component of a sublist. Such chunk mechanisms have also been explored in key-value stores that use a versioned linked list based search structure [5]. We note that this optimization is also applicable to the sublists of DiLi. We do not add this to our implementation to have a fair comparison with skip lists, which cannot have such an optimization across its levels.
The versioned linked list is a classic technique that arose from the need for read-only transactions in databases [20]. Since we only support basic operations insert, remove, and find of a key in this paper, we lack the need for multi-versioning and hence do not compare our performance with versioned lists. Such databases [5, 21] also use lock-free skip lists to boost index search performance. These lock-free skip lists ([28, 11, 10]), can be replaced with DiLi, as it offers comparable performance and horizontal scalability via distribution.
Another well-celebrated direction is to use a universal construction approach, for example, Herlihy [17], to convert sequential implementations to wait-free concurrent implementations. However, universal construction adds overhead compared to direct implementations.
The aforementioned literature are all implementations on a single machine. Although there exist lock-based implementations of a distributed B-Tree in the literature [26], with regard to nonblocking distributed implementations on multiple machines, the literature is rare. The most related work is due to Abe and Yoshiday [2] where a doubly linked distributed list is constructed. The presented strategy uses conflict detection and sequence numbers (with some assumptions) to break symmetries. It guarantees atomic insert and delete operations and non-blocking lookup operations, but the implementation is only obstruction-free.
To the best of our knowledge, DiLi is the first asynchronously distributable conditional lock-free implementation of a linked list, and we believe it is an important step towards constructing other distributed, lock-free, linked data structures (trees, graphs, etc.).
9 Concluding Remarks
In this paper, we focussed on developing DiLi, a distributed linked list data structure. The goal of this data structure is to preserve the client properties of Harris List (specifically, the lock-free nature) while providing horizontal scaling. Specifically, as the size of the data structure grows and as the number of operations performed on it in a given time increases, the data structure needs to be scaled to handle it.
The distribution of the linked list created several challenges. These include ensuring that client operations continue to work correctly in a lock-free manner. Some properties of the list (e.g., linearizability, the number of elements in the list, the number of concurrent operations happening on the sublist) need to be preserved while the list is being split or moved. We developed an innovative approach through offsets of the sublist. This ensured that we can detect a virtual time when no operation is happening on the sublist that is being moved.
Although designed as a linked list, the implementation of DiLi provides a performance comparable to skip lists [11, 28, 10] on a single machine. Additionally, while these skip list implementations cannot provide horizontal scaling via distribution, DiLi is capable of providing such scaling. An implication of this is that any lock-free data structure that is based off of Harris List, can easily be converted into a high-performance distributed data structure using the techniques in this paper.
The data structure considered in this paper, a linked list, differs from other data structures such as hash tables in one important way. The elements in the hash table are not connected to each other. In contrast, elements in a linked list are connected to each other by the notion of the next element. These relations are preserved by the split and move operations in DiLi. It is therefore possible to ensure that these operations can be added to DiLi while ensuring that their lock-free nature is preserved even while the list is being split or moved. This will open up the possibility of horizontally scaling other lock-free data structures (e.g., trees) while preserving their lock-free nature.
Appendix A Registry Operations
398
399Registry *registry;
400 void addEntry(Entry *entry):
[…]
402  do
403     currReg = registry;
404     Registry newReg;
405     for $i=0$; $(i<currReg\rightarrow size\wedge$ $currReg\rightarrow entries[i]\rightarrow keyMin<newKey)$; $i=i+1$ do
[…]
407
[…]
409     for $(;i<currReg\rightarrow size;i=i+1)$ do
[…]
411
412    i = i + 1;
[…]
415
[…]
417Entry *getByKey(Key key):
418  currReg == registry;
419  $\text{entries}=\text{currReg}\rightarrow\text{entries}$;
420
[…]
422  while left $\leq$ right do
423     middle == (left ++ right)//2;
[…]
425        right == middle-1;
426
427    else
[…]
429           left == middle ++1;
430
431       else
432           return entries[middle];
433
434 return null;
435
436
Algorithm 6 Registry Operations
As depicted in Algorithm˜6, the Registry features two main operations – getByKey(key) and addEntry(entry). Add Entry copies the registry to a new registry, while placing the new entry at the right position of the array, that is kept sorted by the keyMin of the entries. In our implementation, we additionally use [25] to safely reclaim old registry (memory stored currReg pointer variable in the addEntry() routine) after the last pointer to it goes out of scope. The operation getByKey() reads the value of the current registry pointer and does the classic binary search on the registry as an array. Note that since we allow multi read (used by client operations), single write (used only by Split) for the registry, any implementation of a lock-based binary search trees that has a wait free search could be used instead of our simple illustration with arrays and copy-on-writes. A removeEntry() method can be implemented by a similar copy-on-write technique.
Appendix B Merge Operation
The algorithm for the Merge operation is depicted in Algorithm˜7. The operation takes the two registry entries of the sublists that are in the same server and next to one another as input, then merges the right sublist with the left sublist and returns the left sublist entry as the response. The algorithm is much similar to Split, but the logic happens in a different order. First, The left sublist entry is updated with the merged key range. Then the right sublist entry is removed from the index. This allows subsequent traversals of the client operations to be possible just from the left sublist. Then the stCt and endCt of the items in the right sublist are updated with the left sublist counters.
Now all that is left before announcing the merge is to remove the subtail and subhead present at the middle of the merged sublist. This is done using a restricted double compare single swap (RDCSS) that is implemented using 3 CAS operations [15]. The operation takes two pairs of <variable, value> pairs (hence the first 4 parameters) for comparison and replaces the first variable with the fifth parameter. For the merge to remove the subtail subhead block, the operation repeatedly tries to update the next pointer of the last element of the left sublist to the first element of the right sublist. To make it asynchronous, it must succeed only when no insertions take place on the subhead to be removed. Hence, the double comparison is performed for a single swap. After this, the new offset is computed the same way as Split (Section˜5.3, by finding the difference between the counters during a momentary write-free period. The updated index is then announced to the other machines via RegisterMergedSublist message.
437Entry *Merge(Entry * leftEntry, Entry rightEntry):
438
[…]
442
443 registry.removeEntry(rightEntry);
444
[…]
446  do
447     prev = curr;
[…]
450     $curr\rightarrownext$;
451
[…]
453 do
[…]
456        $leftLast\rightarrownext$;
457
[…]
460
[…]
462 do
463     $a_{1}$ = $\rightarrowstCt- leftEntry\rightarrowendCt$ ;
464     $a_{2}$ = $\rightarrowstCt- rightEntry\rightarrowendCt$ ;
465
[…]
468
469 for $i\in(serverList-\{me\})$ do
[…]
471
472 return leftEntry;
473
474bool RegisterMergedSublistRecv(Key keyMid, Ref SH):
475  Entry rightEntry = registry.getByKey(keyMid);
476  Entry leftEntry = registry.getByKey(keyMid - 1);
[…]
478  registry.removeEntry(rightEntry);
479  return true;
480
Algorithm 7 Merge Operation
Appendix C Supplementary Results
Figure 4: Scatter plot of the time taken by background operations, plotted against the local elapsed time at which the opertion completed.
Figure˜4 shows the time taken by Split and Move operations (Move includes Switch) at the server that initially owned the entire key range for the experiment. The experiment was a workload of insert operations, that loaded 1M keys into two 2-core machines, with one of them not owning any part of the list key range at the beginning, and begins to receive sublists to own through Move operations the first minute of the load test. We observe that the load of insertions completed at the 180 second mark, after which no Split operations were performed (since all sublists are now within the threshold limit of 125 and the list is no longer receiving updates). After this time, only Moves were performed to complete load balancing.
Appendix D Proof of Correctness for the Background Operation Algorithms
D.1 Properties of stCt and endCt of a Sublist for Delegations
Theorem D.1.
There is only one active subhead pointer to a sublist at any given time.
Proof D.2.
For any sublist, a subhead pointer that stores an item with stCt as negative, would redirect a request either using their newLoc or a fresh registry lookup. However, during a Move, there will exist two subheads with stCt being non-negative – 1. The subhead pointer (SH) of the source machine sublist copy. 2. The subhead pointer (SH*) created in the new machine sublist copy. No registry entry stores SH* during the Move phase. Hence every request gets routed through SH, and will continue to treat it as the active subhead since its stCt is not negative. During a Switch, SH will have its stCt as negative, and thus reroute all requests from the outdated registry entry to the newLoc, which is SH*. Registry entry is then updated to SH*, after which rerouting via SH will no longer be required and SH* will be the only active subhead.
Theorem D.3.
If an update on a sublist has checked the sign of its stCt to be positive after incrementing it, then the stCt of the sublist cannot become negative until the update also increments its endCt.
Proof D.4.
Wen stCt is incremented without incrementing endCt, the difference between the two is more than the offset. This makes it impossible for the computed CAS value check (line 4) to be true until the endCt is also incremented. When the stCt is already set to -$\infty$, no amount of increment can make it positive, thereby making the sublist copy inactive to perform.
D.2 Bound on the Number of Delegations
Theorem D.5.
The maximum number of network round trips involved in a client operation is 2 when there is no Switch operation, and 3 when there is a Switch operation.
Proof D.6.
When there is no Switch operation, the distributed sublist is static. There is exactly one machine that contains the sublist for the corresponding key. Either the client directly requests the corresponding machine (1 total round trip), or requests the machine that forwards this request (2 total round trips). Thus the number of network round trips involved is atmost 2 when there is no Switch operation.
During a Switch operation, a request to the old machine gets delegated to the new machine, bringing about an additional network round trip to the computation. Such a request has to be served in the new machine, as the distributed sublist assumes that another Switch begins long after all existing requests to the old machine for the sublist are completed. Thus, the maximum number of network round trips of a client operation during a Switch is atmost 3.
Since every pair of round trip has a unique thread in common due to the nature of the delegation hops, the number of threads utilized per client operation is one more than the number of network round trips taken for the operation. In other words, the number of threads utilized is 3 when there is no Switch operation and 4 when there is a Switch operation.
D.3 Correctness of the Replay Algorithm in Move
Section˜5.4 shares the intuition behind the replay. The replay algorithm to recreate the moving sublist is done by reimagining the sublist to be built by InsertAfter (prevItem, newItem), instead of Insert(key), as it offers the exact item at which the insertion took place. Note that prevItem and newItem have sId, ts, key members present inside them, for easy notations in the proof. This distinguishes multiple items with the same key that were entered and removed, and ensures that if the prevItem had the same key as an old deleted item, then the replay waits till the newItem is inserted using the logical time stamp. Similarly, Remove(key) is replaced by Delete (prevItem).
Upon receiving a request of the form RepInsert(prevItem, item, oldLocation), the operation is required to find the location of $prevItem$ using their unique (sID, ts) tuple and insert the itemitem at its appropriate place, and return the inserted location back to the requesting server. This server can now set the item matching the $oldLocation$ to have a newLoc{newLoc} provided from the replay message. While finding $prevItem$ and returning a item after insertion is straightforward, performing the insertion at the appropriate place requires some observations. Replaying a Delete is straight forward, as it only needs to find an item of a specific (sID, ts) and ensure that it is marked.
For the sake of this discussion, we say that if $\tt{InsertAfter}$}(Ref(X), Y) has occurred, then Y is a successor of X, and X is a predecessor of Y. We say that X is an ancestor of Y if and only if there is a sequence $X,X_{1},X_{2},\cdots,Y$ such that, each element in the list is a predecessor of the next element. In this case, we also say that $Y$ is a descendent of $X$. For the sake of simplicity, the oldLocation parameter is hidden in the discussion, as it does not affect the replay and is only used to complete assigning a newLoc on callback. We additionally use the terms list item and list node interchangably.
Lemma D.7.
If two requests $\tt{InsertAfter}$} (A,B) and $\tt{InsertAfter}$} (A,C) occur simultaneously, then the node that gets inserted first will have a lower timestamp and will be farther away from A.
Proof D.8.
The lower timestamp clause of the Lemma follows from the fact that the insertion that lost the race in CAS will re-increment the logical clock to get a timestamp higher than the previous values. The latter part of the Lemma follows from the fact that the item from the latter $\tt{InsertAfter}$} will occupy the next pointer of A and will have its own next pointer pointing to the element that was inserted at A earlier.
Lemma D.9.
If there is a replicate message of the form RepInsert(A,C), then $A.ts<C.ts$.
Proof D.10.
This follows from the fact that A must have existed in the sublist before $\tt{InsertAfter}$} (Ref(A), C) was called.
Lemma D.11.
If there is a link of the form $A\rightarrow B$ in the sublist that is being reconstructed, then one out of the following two has to be true: (1) node B was inserted at node A or descendants of A (and so has a timestamp greater than A); (2) B was inserted at a node preceding A (and so has a timestamp less than that of A).
Proof D.12.
This follows from the fact that the entire linked list structure (excluding next pointer mark values) has been created by the use of $\tt{InsertAfter}$} and $\tt Split$} operations. Since $\tt Split$} only inserts special, client-invisible nodes($\tt SubHead$} and $\tt SubTail$}), it is safe to argue that the nodes A and B must have been inserted into the list through $\tt{InsertAfter}$} operations. $\tt{InsertAfter}$} can only insert nodes to the right side of the node provided to it. Hence, if node B is appearing at the right of A, then either it was inserted at A or was inserted at an ancestor of A, at a time before A was inserted into the list. Thus, in the former case, $B.ts>A.ts$ and in the latter case, $B.ts<A.ts$.
Lemma D.13.
If the replicated server has a link of the form $A\rightarrow B$ and $A.ts>B.ts$, then any RepInsert(A,C) must insert C in a place after A, but before B.
Proof D.14.
This follows from Lemma˜D.11, by which B was not inserted at A in this case. Hence B is not a competing insertion at A, and C can be inserted just after A as the first replicated insertion at A.
Lemma D.15.
If the replicated server has a link of the form $A\rightarrow B$, and $A.ts<B.ts$, then any RepInsert(A,C) must traverse from A and find a node D such that $C.ts>D.ts$ and insert C just before D.
Proof D.16.
From Lemma˜D.11, this conveys that B was inserted at A. This means C and B were possibly competing to insert at A, and so if $C.ts<B.ts$, C has to be present after B, according to Lemma˜D.7. After B, there could exist other descendants of A that were inserted in ways similar to B. Hence, the replay has to find the first node D after A, such that $C.ts>D.ts$, which will ensure either that D is not a descendant of A or that D is a descendant of A that was inserted into the list before C was inserted. Either case, C has to be inserted just before D.
Theorem D.17.
The Replay algorithm reconstructs the sublist with the exact structure that is present in the source machine of the $\tt Move$}.
Proof D.18.
Combining Lemma˜D.13 and Lemma˜D.15, the algorithm to replay a request of RepInsert(prevItem,item) condenses down to finding the first node(currcurr in Line 4 of Algorithm˜4), such that $curr.ts<item.ts$ (as in Line 4). This will ensure that currcurr is either a node of type B in Lemma˜D.13 or a node of type D in Lemma˜D.15. Hence, the insert replays reconstruct the same sublist that the corresponding $\tt{InsertAfter}$} operations had created on the other server.
D.4 Termination Condition of the Asynchronous Background Operations
While we will be showcasing the practical termination of the background operations in Section˜7, and proved its asynchronous nature on client operations through lock-freedom, we still identify their termination conditions in this subsection.
Split: While the correctness of Split is straightforward from the explanation in Section˜5.3, the termination of Split operation depends on overcoming three different race conditions – 1. It first competes with the insert operations (Line 3 on $sItem$ to insert the “ST-SH” block into the list. 2. After insertion, it loops to find the new offsets for the split sublists (Line 3). This requires a brief moment in time when no updates take place on either sublist. 3. Finally, it competes with other background Split operations on other sublists to update the registry (Lines 3 to 3).
Move: The correctness of Move comes from the exhaustive proof of correctness of the Replay Algorithm (provided in Section˜D.3) to generate a live clone of the sublist. However, its termination depends on only one race condition - setting stCt to -$\infty$ through the CAS in Line 4 (succeeding this CAS initiates the Switch). This is a type of spin lock, that requires only a brief moment when there are no ongoing updates to the sublist.
Switch: Correctness of Switch comes from the single active subhead pointer property (as in Section˜D.1). However, it however has one race condition – It needs to find the latest subtail of the previous sublist (if one exists) and change its next pointer (Line 5 to 5). This thus competes with Move and Switch operations of that previous sublist, which is practically trivial to overcome as Move of the same sublist should anyway not be invoked too often to avoid performance loss from too many delegations of client operations on this sublist.